which number is a factor of 20?
A) 1
B) 12
C) 40
D) 60

Found 2 solutions by Nate, AnlytcPhil:
Answer by Nate(3500)

What number can evenly go into: 20
20/1 = 20
20/12 = 1.6667
20/40 = 0.5
20/60 = 0.3333
The answer is (A).

Answer by AnlytcPhil(1807)

which number is a factor of 20? 

A) 1

B) 12

C) 40

D) 60

You must learn that the words

"IS A FACTOR OF"

are exactly equivalent to the words

"WILL DIVIDE EVENLY INTO WITHOUT
LEAVING A REMAINDER (OR A REMAINDER
OF 0)."

So pretend the problem were stated this way:

which number WILL DIVIDE EVENLY INTO 20 WITHOUT 
LEAVING A REMAINDER (OTHER THAN 0)? 

A) 1

B) 12

C) 40

D) 60

Let's see. We'll start with the last one.

Try (D).  WILL 60 DIVIDE EVENLY INTO 20 WITHOUT
LEAVING A REMAINDER OTHER THAN 0? Let's see:

    0
60)20
    0
   20

No that leaves a remaider of 20

Try (C).  WILL 40 DIVIDE EVENLY INTO 20 WITHOUT
LEAVING A REMAINDER? Let's see:

    0
40)20
    0
   20

No that also leaves a remaider of 20
   
Try (B).  WILL 12 DIVIDE EVENLY INTO 20 WITHOUT
LEAVING A REMAINDER? Let's see:

    1
12)20
   12
    8

No that leaves a remaider of 8 

Try (A).  WILL 1 DIVIDE EVENLY INTO 20 WITHOUT
LEAVING A REMAINDER OTHER THAN 0? Let's see:

  20
1)20
  2
   0
   0
   0

YUP, that only leaves a remainder of 0,
so that's the answer.

Edwin
AnlytcPhil@aol.com